# Manuel Morante Kristian Frølich Naveed ur Rehman

### v1

@@ 2.2 Multiscale Functional Connectivity
In this study, we propose a novel methodology for analyzing fMRI called Multiscale Functional Connectivity (MFC). The fundamental idea behind this approach is illustrated in Figure 1. Observe that the process consists of a few steps. First, we collected the fMRI data, and we used MMD to unveil IMs associated with that particular individual. Then, we used the multivariate IMs to unveil their corresponding FC patterns. The result is a set of FC patterns that extend across several frequencies, providing a multiscale representation of the FC.
Herein, we explore in detail the most relevant aspects of each step.
Figure 1: Summary of the proposed methodology.
Multivariate Mode Decomposition (MMD)
As we discussed in the introduction, one of the novel points of the proposed approach consists of MMD to extract the natural oscillation within the fMRI data. Formally, MMD is a Signal Processing model that assumes that a multivariate signal of interest accepts a representation as a linear combination of a set of a particular family of amplitude- and frequency-modulated (AM-FM) functions, with a well-defined instantaneous frequency at any given time instance among all the channels (Huang et al., 1998; Rehman and Aftab, 2019). In other words, these intrinsic oscillations, i.e., IM functions, behave similarly to harmonics that remain relatively close to a particular frequency, yet they are flexible enough to accommodate fluctuations in both amplitude and frequency in a data-driven way (Dragomiretskiy and Zosso, 2014).
Unlike other similar alternative studies, such as the approach proposed by (Yuen et al., 2019), MMD inherently exploits the multivariate nature of fMRI data, as Figure 1 illustrates. This is a critical advantage and differentiates it from other approaches. The multivariate nature of MMD allows each ROI to exhibit distinct functional behavior while sharing a common oscillatory behavior, i.e., a common central frequency. This flexibility accommodates a wide range of response changes. This behavior contrasts with conventional approaches, such as ICA, where each voxel/ROI activity is decomposed as a linear combination of a single common time activation.
However, performing MMD constitutes a challenging task, and several methods have been proposed to solve it. Among all the available approaches, in this study, we will focus only on two of the most popular: Multivariate Empirical Mode Decomposition (MEMD) (Rehman and Mandic, 2009) and Multivariate Variational Mode Decomposition (MVMD) (Rehman and Aftab, 2019). Nevertheless, although both methods aim to achieve the same signal decomposition in a fully data-driven manner, they are vastly different from an algorithmic perspective.
On the one hand, MEMD aims to obtain the intrinsic modes using a greedy iterative process. This process unfolds by iteratively averaging the maxima and minima envelopes until we obtain an IM. Although this process is relatively simple for univariate signals (Huang et al., 1998), determining local extrema from multivariate signals is challenging. Rehman and Mandic (2009) solved this problem by projecting the multivariate signal in different directions on hyperspheres. Then, the extrema of the projections are interpolated component-wise to provide the multidimensional envelopes of the signal, which are then averaged to yield the multivariate mean signal. Finally, this average is subtracted from the original signal as an IM function. This process continues iteratively until the residual is sufficiently small.
Observe that this algorithmic procedure causes MEMD to extract the highest frequency first, followed by lower successive frequency components. This characteristic makes MEMD particularly sensitive to noise. Furthermore, due to MEMD’s greedy nature, any mistakes introduced during the extraction of the high-frequency components will spread through the subsequent low-frequency components.
On the other hand, MVMD has the same aim as MEMD but offers better performance (Rehman and Aftab, 2019). Unlike other methods, MVMD aims to estimate all modes simultaneously. The objective is to find an estimate of narrowband functions from the data, where the center frequency of each mode aligns across all the multivariate components of the signal. This renders MVMD more robust to noise and provides a better estimate of the modes.
Functional Connectivity extraction
As illustrated in Figure 1, we can use the obtained multivariate IMs associated with each particular frequency band to uncover the FC at different time scales, providing a complete multiscale FC representation of the fMRI data. Nowadays, we can find a wide range of approaches for estimating the brain’s FC (Bolton et al., 2020b). For simplicity, we propose to use conventional Pearson’s correlation coefficient analysis to calculate the correlation between the time activation patterns within each IM. Note that we can do this because, unlike Yuen et al. (2019) who performed a voxel-wise study, we unveiled a multivariate representation of each IM. Note that this approach differs from static FC analysis, where the raw FC signals are analyzed directly from the ROIs, often after applying static temporal filtering, or from conventional dFC, such as sliding-windows approaches (Lurie et al., 2020).
2.3 Experiments description and fMRI data
In this study, we considered three experiments from the WU-Minn Human Connectome Project (Van Essen et al., 2013). Specifically, we selected the resting state, motor, and gambling experiments from the HCP repository [^1] Human Connectome Project: https://www.humanconnectome.org/. In each experiment, we randomly selected 100 healthy participants aged 22 to 35 years.
The motor experiment followed a standard block paradigm, where a visual cue asked the participants to tap their left or right fingers, squeeze their left or right toes, or move their tongue. Each movement block lasted 12 seconds and was preceded by a 3-second visual cue. Additionally, there were three extra fixation blocks of 15 seconds each. We chose this simple experiment because the activation patterns and neuronal networks involved are well studied (Yeo et al., 2011; Buckner et al., 2011; Turner et al., 2018; Morante et al., 2021), facilitating the evaluation of the results.
Last but not least, the gambling experiment followed a random block paradigm, where participants tried to guess if a randomly generated number between 1 and 9 was either higher or lower than 5. There are two main reasons why we investigated this additional task-related experiment. Firstly, like the motor task, this experiment has been well studied, making it easier to evaluate the results. The second reason is that, unlike the motor task, the gambling experiment’s paradigm is unpredictable, i.e., the guesses of the participants cannot be determined a priori. This randomness adds an unpredictable effect to the responses, increasing the variability in the data and allowing for a more robust and consistent analysis. Additionally, we expect the level of arousal and effort for this experiment to be higher than for the other two experiments, which may appear reflected in the performance of MMD or the obtained inherent FC patterns.
Preprocessing and regions of interest
We obtained the fMRI data directly from the HCP repository [^1] Human Connectome Project: https://www.humanconnectome.org/. The three used datasets were collected using a 3T scanner with a repetition time (TR) of 720 ms. The specific descriptions of the experimental procedures and acquisition parameters are detailed in the HCP imaging protocols [^2] HCP 3T Imaging Protocol Overview: http://protocols.humanconnectome.org/HCP/3T/imaging-protocols.html. Finally, on top of the standard preprocessing pipeline already applied by the HCP (Barch et al., 2013; Van Essen et al., 2013), we further smoothed each brain volume with a 4-mm FWHM Gaussian kernel.
For this study, we divided the brain into several regions of interest (ROIs) using the Automated Anatomical Labeling (AAL) atlas (Tzourio-Mazoyer et al., 2002). Although the AAL atlas maps the entire brain, we only analyzed cerebral regions, which resulted in 90 ROIs.
Table 1: Summary of the selected ROIs organized according to the functional modules described by Parente and Colosimo (2020). The numeric label refers to the 90 ROIs associated with the AAL from (Tzourio-Mazoyer et al., 2002), where odd and even correspond to the right and left hemispheres.
Following the recent module-based network organization proposed by Parente and Colosimo (2020), we grouped these 90 ROIs into seven functional modules. Table 1 contains information regarding the ROIs selected from the AAL and its modules. For the selected 90 ROIs, we extracted the related time series using Nilearn toolbox [^3] Nilearn: https://nilearn.github.io/stable/index.html. Finally, we removed the mean value from each ROI.
Additionally, for the motor task, we added an extra 5 ROIs for the analysis of the time courses associated with the different parts of the motor cortex. For extracting these areas, we used the same motor templates for separating these motor ROIs as, for example, the one implemented by Morante et al. (2021).
3 Results and discussion
We evaluated the proposed methodology in three different fMRI experiments. Firstly, we conducted a comprehensive study of two of the most popular MMD algorithms using fMRI data. Then, we evaluated the neurophysiological relevance of the IMs and their relation to neuronal activity. Finally, we analyzed the FC patterns associated with the FC, their reliability among participants. Additionally, we highlighted the further insights gained through this approach, emphasizing its differences, advantages, and limitations over existing methodologies.
3.1 Intrinsic Mode decomposition analysis and evaluation
The first step of our proposed approach (see Figure 1) consists of extracting the IMs from the fMRI data by performing MMD. In this study, we focused on the two most popular algorithms for this decomposition: MVMD (Rehman and Aftab, 2019) and MEMD (Rehman and Mandic, 2009). Overall, this section aims to illustrate the process of extracting the IMs, understand their physiological meaning relative contributions to the fMRI signal, and, at the same time, examine which algorithm provides better results when working with fMRI data.
MMD results using MVMD
Figure 2: Frequency and energy distribution associated with each mode using MVMD. The boxplot depicts the corresponding results among all the studied participants for the resting-state, motor, and gambling fMRI experiments. The colored area highlights the modes within the neurophysiologically relevant frequency band.
Figure 2 illustrates the frequency (a) and energy (b) distribution of the IMs using the MVMD algorithm for all participants. Firstly, we observed that the first IM exhibits a dominant frequency centered around zero. Similarly, this mode shows a higher relative energy contribution among all the participants. A closer examination of the activation patterns among different individuals revealed that this mode captures trends and low-frequency variations, including residuals and motion-interfering components. We also expected a high energy contribution of this component due to the nature of these low-frequency trends; these components spread over the whole brain, and, therefore, their contribution to the overall detected signal shall be higher.
In addition, although we observed a high degree of consistency in frequency and energy among participants in this mode, a close observation of the actual activation patterns associated with each mode revealed that those trends are very different among participants, which indicates that those low-frequency trends are individual-dependent.
Modes 2 to 5, highlighted within a shaded area in Figure 2, appeared within the neurophysiologically relevant frequency range. This range includes typical brain activity frequencies related to several cognitive and neurophysiological processes (Cordes et al., 2001). Table 2 shows the average central frequencies for each IM among participants, and we found that those results were relatively consistent among participants and experiments.
Resting state Motor Gambling Freq. BW (mHz) Freq. BW (mHz) Freq. BW (mHz) Mode 1 0 6.8$\pm$ 1.0 0 21.6$\pm$ 1.5 0 26.3$\pm$ 1.5 Mode 2 23 14.3$\pm$ 1.9 23 36.5$\pm$ 4.1 26 44.9$\pm$ 4.3 Mode 3 54 15.0$\pm$ 2.2 56 40.5$\pm$ 1.9 66 50.8$\pm$ 2.3 Mode 4 96 15.9$\pm$ 2.7 99 43.0$\pm$ 3.1 115 53.0$\pm$ 2.9 Mode 5 156 16.9$\pm$ 2.7 157 47.6$\pm$ 3.9 182 57.4$\pm$ 3.6 Mode 6 232 16.5$\pm$ 2.8 231 50.6$\pm$ 3.6 257 57.0$\pm$ 3.8 Mode 7 310 17.2$\pm$ 3.1 308 50.8$\pm$ 3.9 330 57.6$\pm$ 4.1 Mode 8 407 18.1$\pm$ 2.5 400 54.1$\pm$ 4.5 419 61.6$\pm$ 4.4 Mode 9 521 18.3$\pm$ 2.4 514 56.6$\pm$ 3.9 533 63.8$\pm$ 4.0 Mode 10 632 17.4$\pm$ 2.3 627 53.8$\pm$ 5.2 627 56.8$\pm$ 7.4
Table 2: Average frequency and their corresponding bandwidth (BW) associated with each IM for the different studied experiments for MVMD.
The remaining modes spread across high frequencies. As discussed in Section 2.1, signals within this frequency range originate from a mixture of different interfering components. These components include those induced by respiration movements, heartbeat, and cerebrospinal fluid pulsations. Notably, mode 6 –approximately centered at 250 mHz– captures the primary respiratory-related component, while mode 9, with a central frequency of approximately 520 mHz, aligns well with the first harmonic of the respiratory frequency (Cordes et al., 2001; Yuen et al., 2019).
Regarding the relative energy contribution, according to Figure 3 (b), which illustrates the relative energies of IMs among all participants. We observed that, in general, modes with lower frequencies exhibited higher energy, while those with higher frequencies had less and less energy contributions. Specifically, the first three modes contain the most energy, whereas the high-frequency modes contribute comparatively little.
Overall, we observed that the results obtained from MVMD align with the expected behavior of the data as described in the existing literature, as we have briefly summarized in Section 2.1. These results validate MVMD as a highly suitable method for this particular task, as it effectively captures the different IMs naturally present in fMRI among participants.
MMD results using MEMD
Figure 3: Frequency and energy distribution associated with each mode using MEMD. The boxplot depicts the corresponding results among all the studied participants for the resting-state, motor, and gambling fMRI experiments.
Figure 3 illustrates the frequency (a) and energy (b) distribution of the modes derived from analysis using MEMD among all the participants. However, in contrast to the results of MVMD shown in Figure 2, the results from MEMD exhibit an entirely different behavior.
First, we observed that the number of intrinsic modes recovered by MEMD varies among individuals and experiments. Figure 3 shows that the maximum number of IMs among experiments is also different per experiment. No doubt, variations in the number of IMs per participant and experiment are expected; it is natural to assume that not all individuals activate the exact number of IMs, particularly if they are performing different tasks. In this regard, MEMD offers more flexibility by using a greedy approach and, unlike MVMD, which fixes the number of IMs to a specific value, MEMD finds all potential components without the necessity of setting any prior guess. However, although this feature is desirable in practice, the greedy approach makes MEMD particularly sensitive to noise. In this way, the results obtained from MEMD are poor and seem rather unusual due to several factors we discuss below.
On the one hand, we observed that the resting-state experiment was the experiment that exhibited the maximum number of IMs, displaying at most 11 modes. In contrast, gambling and motor only exhibited a maximum of 10 and 9, respectively. This unexpected result contradicts the expected nature of these experiments; in the resting state, participants were at rest and calm. Consequently, we would have expected to find the minimum number of components when the participants rested. On the other hand, participants were more actively engaged during the performance of the gambling experiments. Therefore, we would have expected to find the maximum number of active IMs associated with this experiment instead.
Similarly, the overall frequency distribution is odd. We found that the modes appeared predominantly in the low-frequency range. For instance, all the modes, except the last two from the resting-state experiment, are below 100 mHz. Table 3 shows the central frequencies and bandwidths associated with these modes. From observing this table, it is evident that many modes appeared to be dominated by the low frequencies and many of those modes show overlapping bandwidths.
Resting state Motor Gambling Freq. BW (mHz) Freq. BW (mHz) Freq. BW (mHz) Mode 1 1 4.6$\pm$ 0.1 5 19.4$\pm$ 0.3 5 24.1$\pm$ 0.7 Mode 2 2 5.1$\pm$ 0.8 6 20.6$\pm$ 4.6 9 27.3$\pm$ 5.4 Mode 3 3 6.7$\pm$ 1.2 12 27.6$\pm$ 5.0 18 37.4$\pm$ 7.3 Mode 4 6 8.5$\pm$ 1.1 23 35.4$\pm$ 4.3 34 46.5$\pm$ 4.6 Mode 5 11 9.9$\pm$ 1.2 44 41.6$\pm$ 5.4 63 54.4$\pm$ 4.9 Mode 6 19 12.1$\pm$ 1.9 81 51.9$\pm$ 8.3 109 68.2$\pm$ 10.2 Mode 7 32 14.5$\pm$ 2.4 142 60.6$\pm$ 9.3 202 78.3$\pm$ 12.7 Mode 8 51 14.4$\pm$ 3.3 192 60.2$\pm$ 10.5 222 80.2$\pm$ 13.3 Mode 9 82 12.8$\pm$ 3.4 190 62.0$\pm$ 5.4 244 76.2$\pm$ 19.2 Mode 10 140 12.6$\pm$ 2.5 170 52.8$\pm$ 9.6 Mode 11 153 12.3$\pm$ 2.5
Table 3: Average frequency and their corresponding bandwidth (BW) associated with each IM for the different studied experiments for MEMD.
We were surprised by not observing any high-frequency component related to respiratory or cardiac pulsations. High-frequency contributions from physiological sources, including respiratory components, as well as from cerebrospinal fluid and cardiac pulsations –as we discussed in previous sections– constitute a natural part of the fMRI signal. Somehow, these contributions are absent. Overall those results contrast strongly with the results from MVMD in Figure 2, which was able to capture the full range of frequencies naturally associated with fMRI data.
Finally, when analyzing the energy distribution in Figure 3, we observed that the first mode in all the experiments exhibited the most energy contribution, as expected due to the nature of the first mode, which is in line with the discussion of the MVMD results. However, unlike the results in Figure 2, we observed that the energy associated with the rest of the IMs from MEMD exhibited almost a homogeneous distribution.
We hypothesized that MEMD led to those unnatural and unreliable results due to the combination of two factors. First, the nature of the noise in the data, fMRI data is corrupted with Rician noise Poldrack et al. (2011), with especially high levels of energy at low frequencies due to motion artifacts. Second, the greedy nature of MEMD means that the algorithm is mostly driven by the noise associated with the first retrieved components rather than the inherent oscillation of the data.
In addition, another potential problem that might have influenced the results of MEMD is the multivariate extension of this algorithm. As illustrated by Yuen et al. (2019), the univariate version of these algorithms, namely EMD and VMD, exhibited similar results. However, this is not the case we observed for the multivariate case. On the contrary, we observed that the results obtained from MEMD exhibit an entirely different behavior.
Overall, we observed that MEMD seems to be driven by low-frequency trends, as most of the components appear dominated by those low trends (see Table 3). Therefore, after all these observations, we conclude that MEMD appears unsuitable for MMD in fMRI data, as it is unable to extract reliable information from the data.
Frequency and energy differences between fMRI experiments
Based on the results from the previous section, it is clear that MVMD is an excellent candidate for MMD in fMRI data. Specifically, it provides a good separation of the intrinsic modes of the fMRI across the expected natural frequencies. This section focuses on the analysis of the results obtained from MVMD among the studied fMRI experiments.
Overall, the frequency and energy distribution for MVMD in Figure 2 show a similar trend among all the studied experiments. Nonetheless, upon closer examination, we noticed some interesting differences among experiments.
For instance, modes 5 and 6 in the gambling experiment exhibited a significantly higher central frequency than the other experiments. Similarly, when examining the energy distribution of these modes, we observe that modes 6 and 7 significantly contribute more to the overall energy of the signal.
The nature of mode 6 makes this result particularly interesting. These findings suggest that the participant’s physiological state during the gambling experiment may have been different from the resting-state and the motor task experiments. The higher frequency rate of the respiratory component associated with mode 6, as well as the high energy contributions in modes 6 and 7, which contain cardiac contributions, seem to indicate faster physiological rhythms. This particular physiological state corresponds with a higher level of arousal, which potentially reflects a state of excitement or nervousness from the gambling experiment, which contrasts with the potential physiological states from the other experiments.
Moreover, a closer examination of the energy distribution of low-frequency intrinsic modes among the different experiments reveals a significant difference in their energy contributions. In the resting-state experiment, the majority of energy is concentrated in the first mode, representing the dominant and most energetic component. This mode is followed by modes 2 and 3, which also exhibit significant energy levels. In contrast, in the task-related experiments, the first mode is less dominant, and neurophysiologically relevant modes (modes 2-5) consistently show higher energy levels. Modes 2 and 3 capture a substantial portion of the energy, while modes 4 and 5, although showing lower energy contributions, are higher compared to the resting state.
The difference in energy contributions of neurophysiological IMs between resting-state and task-related experiments suggests that the brain is more engaged and active during task performance compared to a state of rest, which makes intuitive sense. The higher energy levels in neurophysiologically associated modes indicate increased neural activity and processing during the task, highlighting the cognitive demands and involvement required for task completion.
3.2 Further analysis of the nature of the IMs
As Figure 1 illustrates, the MMD decomposition offers an alternative way to decompose the fMRI data in inherent modes. As illustrated in Figure 1, each IM provides a multivariate response where each ROI exhibits its particular response. However, despite this flexibility, all of those time activation patterns are characterized by sharing a common oscillatory behavior (Rehman and Aftab, 2019). Therefore, having such flexibility, it is still unclear how those different time activation patterns behave among ROIs, and it is unclear if they still provide reliable and consistent results among participants.
Consequently, to further shed light on these questions, we focus our attention only on the neurophysiologically relevant modes, i.e., IMs 2, 3, 4, and 5. For this study, we decided to investigate only the time activations using the task-related experiments because, unlike the resting-state experiment, for the task-related experiments, we have access to experimental paradigms, which allows us to approximately define the actual hemodynamic response of the brain and use it for comparisons.
Examination of the activation patterns for the motor task experiment
For the motor task experiment, we selected the primary visual cortex (ROI 43 and 44) for the visual responses. In addition, for this study, we further divided the ROI from the motor cortex into five additional motor-related ROIs for the different motor areas related to the right/left hands, feet and tongue. Finally, we performed MMD again with these additional ROIs.
Figure 4 illustrates the average time courses for the relevant ROIs associated with the motor experiment among all the participants. The reported lines correspond to the average time activations for modes 2 and 3 among all the studied participants (without any additional postprocessing). Finally, the orange lines in Figure 4 represent the canonical task-related component expected within each main ROI. We obtained the task-related components using the classical convolutional model with the canonical Hemodynamic Response Function (HRF) (Power et al., 2011). This visualization is crucial because it reflects MVMD’s capability to reveal neurophysiological information, and how each mode contributes to the brain activity.
Figure 4: Average time courses among all the studied participants associated with the first neurophysiological modes (2 and 3) from the results of MVMD (blue lines), and the canonical task-related component (orange line) for the specified ROIs associated with the different motor tasks.
Interestingly, although mode 3 seems better suited to visual cues, we can see that it also contributes to the motor cortex, which indicates a potential connection between the visual and motor-related areas. This is a very interesting feature because this allows us to we can explore how the different areas interact at different timescales.
Overall, these findings provide strong evidence that MMD is an effective method for analyzing brain activity and extracting meaningful information from different regions of interest. The fact that mode 2 aligns closely with block-related activity in the motor cortex ROIs, while mode 3 captures the visual cues associated with the motor task and fixation, demonstrates the interpretability and naturalness of the results obtained through MMD. This further supports the validity and usefulness of MMD in studying the connection between different brain regions and their functional roles.
3.3 Multiscale Functional Connectivity
Reliability of the FC maps among participants
Reliability refers to the ability of methods to consistently detect significant activity within the expected ROIs (Morante et al., 2020). In this case, we focus the attention on the reliability associated with the FC patterns from MMD among participants. Since all participants are analyzed independently, failing to provide consistent results among participants indicates that the method is unable to capture relevant common information.
Therefore, we conducted a study to assess the reliability of the FC among all participants. This study aimed to demonstrate how different FC patterns behaved across participants. In particular, we studied the individual reliability of the static FC patterns associated with each IM for the three considered fMRI experiments.
For completeness, we evaluated the FC across all IMs. To this end, we calculated the Pearson’s correlation obtained from all the possible pairs of comparisons across all participants. We want to emphasize that this step was critical in ensuring the validity and generalizability of our findings, as it allowed us to confirm that the FC patterns associated with each IM were consistent and reliable across all participants, providing a solid foundation for our subsequent analyses.
@@ 3.3 Multiscale Functional Connectivity
Figure 5: Individual reliability of the static FC patterns associated with each intrinsic mode for the three considered fMRI experiments using MVMD. The boxplots depict the Pearson’s correlation values obtained among all the possible individual pair comparisons across all the participants
Upon examination, we observed that all experiments followed a similar trend. Overall, modes within the neurophysiological activity band exhibited higher similarity than the other modes. Specifically, modes 3 and 4 demonstrate remarkable reliability.
In contrast, mode 1 exhibited the lowest similarity across participants in all experiments. These results indicate that the observed patterns associated with the first mode largely depend on each individual; as we already pointed out in the previous section, the first mode contains individual trends and motion residuals that depend on each particular individual.
Regarding the remaining modes above the neurophysiological band, even though they are less consistent than modes within the neurophysiological band, they still exhibit a higher similarity to mode 1. This result indicates that the FC associated with these IMs exhibits a certain degree of consistency among individuals, despite mostly containing interfering physiological components as previously discussed.
Although the three studied experiments exhibited a similar trend, a closer examination reveals some relevant differences. First, the resting-state experiment consistently showed increased similarity for all modes, with mode 4 (96 mHz) being the most consistent among all individuals. This may also be related to why resting-state is the dominant approach in functional connectivity (FC) studies; on top of its simplicity, these results highlight that, in contrast to task-related experiments, activation patterns at different frequencies remain stable and produce more reliable FC patterns among many individuals.
Interestingly, if we focus on the high-frequency modes, modes 8 and 9 maintain a relatively higher similarity than modes 7 or 6. We speculate that this is due to cardiac interfering components at rest, which provide similar patterns. Cardiovascular and cerebrospinal pulsations during rest are more stable and predictable, leading to consistent connectivity patterns across individuals in the high-frequency modes.
Conversely, for the motor task, we observed that mode 3 (56 mHz) was the most consistent among all the participants, followed closely by mode 4. This result was expected since mode 3 (56 mHz) contains a visual cue common to all participants. The presence of a shared stimulus in this mode likely contributes to its higher consistency among individuals during the motor task.
@@
For completeness, we further evaluated the reliability of the modes obtained using the MEMD algorithm. Figure 6 illustrates the obtained results using MEMD.
Figure 6: Individual reliability of the static FC patterns associated with each intrinsic mode for the three considered fMRI experiments using MEMD. The boxplots depict the Pearson’s correlation values obtained among all the possible individual pair-comparisons across all the participants
No doubt, we observed that MEMD results exhibited an entirely different behavior. In contrast to the results from Figure 5, we noted that the first modes of MEMD are poorly similar among participants, whereas the high-frequency modes exhibited the highest similarity. As we previously hypothesized, these results support the idea that the MEMD algorithm seems to be mostly driven by the noise associated with the first retrieved components rather than the inherent oscillation of the data.
Overall, we can see that those results are difficult to interpret and are inconsistent with the expected behavior of the data. Therefore, these results further evidence that MEMD is unsuitable.
MFC patterns and analysis
After understanding the frequency and energy distribution of the different IMs, as well as their reliability among individuals, we focus on the analysis of the FC patterns associated with the IMs. In this case, we are interested in studying only neurophysiological activation patterns. Therefore, we analyzed only the FC patterns associated with modes 2, 3, 4, and 5.
Figure 7 depicts the average FC patterns for each mode. We obtained those FC patterns by averaging the individual FC patterns across all participants. Each row corresponds to a particular mode, and each column contains different experiments. For all comparisons, we performed a statistical test with respect to a null dataset generated from each particular decomposition by randomly mixing the temporal samples of the IMs. Pearson’s correlation coefficients were Fisher-Z transformed. The lower diagonal of each connectivity matrix displays the average correlation coefficients, while the upper diagonal shows only the correlation values that also exhibited significant activation compared to the null data derived from a permutation-based t-test corrected to false discovery rate adjusted to $p<0.001$. For convenience, we arranged the ROIs according to the leading module (left and right), following the order reported in Table 1.

### v2

@@ 4.5 Multiscale Functional Connectivity
Once the IMs have been extracted and identified with neurophysiological information, their FC can be extracted. Nevertheless, we examined the reproducibility of participants’ results to ensure relevance. We need to ensure that the IM actually produces consistent results among participants, as extraction of the IM is performed individually. This reproducibility study enabled us to confirm the validity of our proposed approach.
In this regard, note that this step differs from other approaches, such as [43], in that it constitutes a necessary step to ensure results consistency. Rather, our approach is robust enough to be shown after IM extraction, without any fine-tuning.
Reproducibility of the FC maps among participants
Reproducibility refers to the ability of methods to consistently detect consistent activity within the expected ROIs among different experimental realizations [44, 24]. In this case, we focus on the reproducibility associated with the FC patterns from MMD among participants. Since all participants are analyzed independently, failing to provide consistent results among participants indicates that the method is unable to capture relevant common information.
Therefore, we conducted a study to assess the reproducibility of the FC among all participants. This study aimed to demonstrate how different FC patterns behaved across participants. In particular, we studied the individual reproducibility of the static FC patterns associated with each IM for the three considered fMRI experiments.
For completeness, we evaluated the reproducibility of the FC patterns across all IMs. To this end, we calculated the Pearson’s correlation obtained from all the possible pairs of comparisons across all participants. We want to emphasize that this step was critical in ensuring the validity of our findings, as it allowed us to confirm that the FC patterns associated with each IM were consistent across all participants, providing a solid foundation for our subsequent analyses.
@@
Figure 8: Reproducibility of the static FC patterns associated with each intrinsic mode for the three considered fMRI experiments using MVMD. The boxplots depict the Pearson’s correlation values obtained among all the possible individual pair comparisons across all the participants
Upon examination, we observed that all experiments followed a similar trend. Modes within the neurophysiological band exhibited higher similarity than the other modes. Specifically, modes 3 and 4 demonstrate remarkable reproducibility.
In contrast, mode 1 exhibited the lowest similarity across participants in all experiments. These results indicate that the observed patterns associated with the first mode largely depend on each individual; as we already pointed out, the first mode contains individual trends and motion residuals that depend on each particular individual.
Regarding the remaining modes above the neurophysiological band, even though they are less consistent than modes within the neurophysiological band, they still exhibit a higher similarity to mode 1. This result indicates that the FC associated with these IMs exhibits a certain degree of consistency among individuals despite mostly containing interfering physiological components, as previously discussed.
Although the three studied experiments exhibited a similar trend, a closer examination reveals some interesting differences. First, the resting-state experiment consistently showed increased similarity for all modes, with mode 4 (96 mHz) being the most consistent among all individuals. This may also be related to why resting-state is the dominant approach in FC studies; on top of its simplicity, these results highlight that, in contrast to task-related experiments, activation patterns at different frequencies remain stable, as they seem to produce more reliable FC patterns.
If we focus on the high-frequency modes, modes 8 and 9 maintained a relatively higher similarity than modes 7 or 6. We speculate that this is due to cardiac interfering components at rest. Cardiovascular and cerebrospinal pulsations during rest are more stable and predictable, leading to consistent correlation patterns across individuals in the high-frequency modes [22].
Conversely, for the motor experiment, we observed that mode 3 (56 mHz) was the most consistent among all the participants, followed closely by mode 4. This result was expected since mode 3 (56 mHz) contains a visual cue common to all participants. The presence of a shared stimulus in this mode likely contributes to its higher consistency among individuals during the motor task. This may also indicate that the visual cue dominates the brain activity among all the participants.
@@
Further evaluation of MEMD
Although the results in the previous section already evidence the inadecuacy of MEMD to extract relevant information from fMRI data, we decided, for completeness, to further evaluated the reproducibility of the modes obtained using the MEMD algorithm. Figure 9 illustrates the obtained results using MEMD.
Figure 9: Reproducibility of the static FC patterns associated with each intrinsic mode for the three considered fMRI experiments using MEMD. The boxplots depict the Pearson’s correlation values obtained among all the possible individual pair-comparisons across all the participants
As expected, the FC associated ot the IMs from MEMD are not similar among participants. Interestingly, the high-frequency modes exhibited the highest similarity. These results further support the idea that the MEMD algorithm seems to be mostly driven by the noise associated with the high-frequency components rather than the inherent oscillation of the data.
4.6 Neurophysiological MFC patterns and analysis
After understanding the frequency and energy distribution of the different IMs, their corresponding FC patterns and the reproducibility among individuals, we focus on the interpretation of the FC patterns associated with the IMs. In this case, we are interested in studying only neurophysiological activation patterns. Similarly, we focus our study on the neurophysiological FC patterns from MVMD, as it provided the best performance.
Figure 10 depicts the average FC patterns for each mode. We obtained those FC patterns by averaging them across all participants, as they exhibited a high reproducibility. Each row corresponds to a particular mode, and each column contains different experiments. For all comparisons, we performed a statistical test with respect to a null dataset generated from each particular decomposition by randomly mixing the temporal samples of the IMs. Pearson’s correlation coefficients were Fisher-Z transformed. The lower diagonal of each connectivity matrix displays the average correlation coefficients, while the upper diagonal shows only the correlation values that also exhibited significant activation compared to the null data derived from a permutation-based t-test corrected for false discovery rate adjusted to $p<0.001$. For convenience, we arranged the ROIs according to the leading module (left and right), following the order reported in Table 1.